# Silent Speech Sentence Recognition with Six-Axis Accelerometers using Conformer and CTC Algorithm

### v2

@@ 4 Discussion
The strength of the proposed method lies in its ability to dynamically recognize arbitrary silent sentences by autonomous segmentation and identification based on a word database, without prior knowledge of the sentences. This overcomes the limitations of previous studies that treated word classification or treated sentences as a whole entirety similar to the word [22, 45]. The previous methods are only capable of static recognition and lack flexibility and expressiveness, while this approach allows for a more versatile and practical SSI system.
@@
In addition to accuracy and model performance, the practical deployment of silent speech interfaces must consider sensor comfort. Future iterations could explore more ergonomic designs, such as thinner flexible substrates or skin-conformal electronics, to enhance wearability over extended periods.
Furthermore, continuous collection and processing of biometric motion data may raise privacy concerns. Ensuring data security, anonymization, and user control over data access will be essential for the responsible adoption of such systems in real-world healthcare or communication settings.
The proposed method offers a promising solution for individuals with speech and communication impairments, paving the way for more practical and accessible SSI systems. Future research would focus on addressing the challenges mentioned above and exploring advancements to achieve better accuracy and generalization.
@@ 5 Conclusion
In this paper, a new method has been presented for silent sentence recognition. The system focuses not only on word-level expression but also on complete sentence (sequence) output. Accelerometers combined with a Conformer-based neural network was applied to extract the facial motions and then translate mechanical motions into word and sentence transcriptions accurately. The average accuracy in sentence recognition and word classification is 97.17%. The proposed method performs well on easily confused word recognition and shows robustness and generalization ability on the few-shot task.
This work is the first to recognize silent sentences dynamically and promotes the practical implementation and utilization of the system. The system is portable and applicable for patients in their daily usage. In the future, the integration of accelerometers and other modalities, such as microphones, will be explored.
Acknowledgments

### v3

@@ 4 Discussion
The strength of the proposed method lies in its ability to dynamically recognize arbitrary silent sentences by autonomous segmentation and identification. This overcomes the limitations of previous studies that treated word classification or treated sentences as a whole entirety [14, 15]. The previous methods are only capable of static recognition and lack flexibility and expressiveness, while this approach allows for a more versatile and practical SSI system.
@@
In addition to model performance, the practical deployment of silent speech interfaces must consider sensor comfort. Future iterations could explore more ergonomic designs, such as thinner flexible substrates or skin-conformal electronics, to enhance wearability over extended periods.
The proposed method offers a promising solution for individuals with speech and communication impairments, paving the way for more practical SSI systems. Future research would focus on addressing the challenges mentioned above and exploring to achieve better accuracy and generalization.
@@ 5 Conclusion
This paper presents a silent sentence recognition system using accelerometers and a Conformer-based network to capture facial motions and translate them into word and sentence transcriptions. The model achieves an average accuracy of 97.17% at both word and sentence levels, demonstrating strong performance on classification and generalization.
As the first dynamic silent sentence recognition approach, the system offers portability and practical usability for patients in daily scenarios. Future work will explore multi-modal integration with microphones and other sensors.
6 Acknowledgement